OpenStudy (anonymous):

simplify the expression: cosx(cosx-secx) should i make sec look like sin & cos first?

OpenStudy (anonymous):

Well, secx = 1/cosx. We then have: cosx(cosx-(1/cosx)) = cos^2x - 1 = -sin^2x, because of the trig identity that sin^2x + cos^2x = 1 :-)

OpenStudy (anonymous):

So yes, as stated on your other question about simplifying trig equations, always break up sec, tan, cossec, cotan, etc.

OpenStudy (anonymous):

thx ^^

OpenStudy (anonymous):

Distribute. You will get: \[\cos ^{2}(x) - \cos(x)\sec(x)\] \[\cos ^{2}(x) -\frac{\cos(x) * 1}{\cos(x)}\] \[\cos ^{2}(x) - 1\] \[-\sin ^{2}(x)\]

OpenStudy (anonymous):

thx again ^^

OpenStudy (anonymous):

X)